OpenStudy (anonymous):

A couple decided to have 5 children. (a) What is the probability that they will have at least two girls? (b) What is the probability that all the children will be of the same gender?

OpenStudy (anonymous):

what is the probability of a baby being a boy or girl?

OpenStudy (anonymous):

1/2 right?

OpenStudy (anonymous):

similar to the problem if they decide to have 3 kids with two girls and one boy {BBB,BBG,BGB,BGG**,GBB,GBG** ,GGB**,GGG} => so probability of two Girls(G) ) and oneBoy(B) P(2G union) 1B) = 3/8 ..do similar way for 5 kids and try to understand the answer.

OpenStudy (anonymous):

Use binomial theorem to get the answer. ( take (x+y)^3 x==> represents girls and y==> boys. (x+y)^3 =1 x^3 + 3x^2y + 3xy^2 + 1y^3 the coefficient of the term that represents x^2y(2 girls and one boy) has 3 and total coefficient sum=8 so probability=3/8 The binomial theorem can be used with following characteristics * There are only two possibilities ...success or failure * The events are independent * The probability of each outcome is same for each trial Such problems are called Bernauli trials . Let p represents the probability of success. let q represents the probability of failure. nCrp^(n-r)q^r represents the probability of exactly r successes. Use that here 5C2xy^2 [ 2 girls and one boy] or use tht to expand to get the coefficients use that for (x+y)^5 fr 5 children = x^4 +

OpenStudy (anonymous):

I got the right answer for B but can't get the right answer for A...

OpenStudy (anonymous):

(x+y)^5 = x^5 + 5x^4y + 10x^3y^2 + 5xy^4 + 10x^2y^3 + y^5 let x => represent girl and y=> represent boy. In 5 kids, total same gender (either all boys(y) or all girls(x) are the term x^5 or y^5 = 1 out of 32 =1/32

OpenStudy (anonymous):

1/32 is not the answer for part A

OpenStudy (anonymous):

2nd part all boys or all gilrs

OpenStudy (anonymous):

GGGGG GGGGB GGGBB GGBBB GBBBB if we work out the probability that there is only 1 girl, and take that away from 1, we have the probability that there is at least 2 girls

OpenStudy (anonymous):

probabiliy of two girls(x) = 10x^2y^3 , so the coefficient of this term is 10 ==> probability =10/32

OpenStudy (anonymous):

10/32 is not the answer for A

OpenStudy (anonymous):

I already have the answer for B so you don't need to solve that. But A is wrong

OpenStudy (anonymous):

GBBBB - 1/32 BGBBB - 1/32 BBGBB - 1/32 BBBGB - 1/32 BBBBG - 1/32 total = 5/32 answer = 1- 5/32 = 27/32

OpenStudy (anonymous):

27/32 is wrong too...

OpenStudy (anonymous):

25/32

OpenStudy (anonymous):

im pretty sure i did it right... unless p =/= 1/2

OpenStudy (anonymous):

I did it the same way as you and got the same answer but it's wrong.

OpenStudy (anonymous):

@experimentX maybe you can clear this up

OpenStudy (experimentx):

lol ... i am always prone to mistake on first try.

OpenStudy (anonymous):

:) im going out in a bit so i thought i should leave him in capable hands, if you're willing to help him

OpenStudy (anonymous):

cya guys!

OpenStudy (experimentx):

yeah, i will try my best.

OpenStudy (anonymous):

Haha thank you :)

OpenStudy (anonymous):

5x^4y + 10x^3y^2 + 10 x^2y^3 + x^5(this term all girls) a). prob tht they will will hve at least 2 girls = 5+10+10+1/32 =26/32

OpenStudy (anonymous):

Wow. 26/32 is actually right. Can't believe I was off by only one number.

OpenStudy (anonymous):

Thank you all so much! Appreciate it

OpenStudy (experimentx):

i kinda think like @rumadutt

OpenStudy (anonymous):

z11 you are doing without understanding. try to understand it.

OpenStudy (anonymous):

use binomial theorem if it satisfies those conditions

OpenStudy (experimentx):

seems like binomial probability or so what it is called i forgot ...

OpenStudy (anonymous):

Ok I get it. THANKS

OpenStudy (experimentx):

i guess this is it http://en.wikipedia.org/wiki/Binomial_probability